# Pseudo-Asynchronous Local SGD: Robust and Efficient Data-Parallel Training

### v1

@@ 8 Discussion and Conclusion
  • We empirically validated PALSGD on image classification and language modeling tasks, demonstrating its effectiveness in reducing training time and improving model performance compared to baseline methods such as DDP, Local SGD and DiLoCo (Section 7, Figures 1, 3, and 2).
@@ A.1 Assumptions and Main Result
$\mathbb{E}[F(\hat{x}_{T})]-F(x^{*})\leq\tilde{O}\Bigg{(}\frac{LHR_{0}^{2}}{p}\cdot\exp(-\frac{pT}{\kappa H})+\frac{\sigma^{2}}{\mu KT}+\frac{(\kappa+p^{-1}H)\sigma^{2}}{\mu T^{2}}\Bigg{)}.$
@@ A.2 Basic Definitions
For any $k\in[K],t=0,\cdots,T-1$, let $b_{k}^{(t)}$ denote a Bernoulli random variable with parameter $p$ (i.e., $b_{k}^{(t)}=1$ with probability $p$ and $b_{k}^{(t)}=0$ with probability $1-p$). Let
@@ A.3 Main Technical Lemmas
[…]
@@ A.4 Proof of Theorem 2
[…]
@@
Next, by taking the weighted average from $t=0$ to $T-1$ with weight $w_{t}=(1-\mu\alpha)^{-(t+1)}$ and normalizing factor $W_{T}=\sum_{t=0}^{t-1}w_{t}$, we have
@@
[…]
@@
[…]
@@
[…]
@@
If $\frac{p}{48LH}\leq\frac{\ln(\mu^{2}d_{0}T^{2}K/\sigma^{2})}{\mu T}$, then
$\frac{1}{W_{T}}\sum_{t=0}^{T-1}w_{t}e_{t}\leq\tilde{O}\left(\frac{LHd_{0}}{p}\exp(-\frac{\mu pT}{LH})+\frac{LH^{2}\sigma^{2}}{\mu^{2}T^{2}}+\frac{\sigma^{2}}{\mu KT}\right)$
Otherwise, if $\frac{\ln(\mu^{2}d_{0}T^{2}K/\sigma^{2})}{\mu T}<\frac{p}{48LH}$, then
$\frac{1}{W_{T}}\sum_{t=0}^{T-1}w_{t}e_{t}\leq\tilde{O}\left(\frac{\sigma^{2}}{\mu TK}+\frac{LH^{2}\sigma^{2}}{\mu^{2}T^{2}}\right)$
Therefore, assuming $p\leq\frac{1}{\kappa H}$ we have
$\frac{1}{W_{T}}\sum_{t=0}^{T-1}w_{t}e_{t}$ $\leq\tilde{O}\left(\frac{LHd_{0}}{p}\exp(-\frac{\mu pT}{LH})+\frac{LH^{2}\sigma^{2}}{\mu^{2}T^{2}}+\frac{\sigma^{2}}{\mu KT}\right)$
$\leq\tilde{O}\left(\frac{LHd_{0}}{p}\exp(-\frac{\mu pT}{LH})+\frac{(L+\mu p^{-1}H)\sigma^{2}}{\mu^{2}T^{2}}+\frac{\sigma^{2}}{\mu KT}\right).$
@@ A.5.1 Proof of Lemma 2
where $b_{k}^{(t)}$ is a Bernoulli random variable that equals to $1$ with probability $p$ and 00 with probability $1-p$. Also $\xi_{k}^{(t)}$ denotes the data sampled by the $k$-th server at step $t$. Let $\mathcal{F}_{t}=\{\xi_{k}^{(s)}\}_{s=0,\cdots,t-1,k\in[K]}\cup\{b_{k}^{(s)}\}_{s=0,\cdots,t-1,k\in[K]}$. Conditional on any fixed $\mathcal{F}_{t}$, we then rewrite each of the summands as
@@
[…]
@@ A.5.2 Proof of Lemma 3
[…]
@@ A.5.3 Proof of Lemma 4
[…]
@@
[…]
@@
[…]
@@ Appendix B Experimental Details
We conducted our experiments on four workloads: CIFAR-10, ImageNet-1K, TinyStories and Wikitext - QNLI. The CIFAR-10 and ImageNet-1K datasets were used for image classification tasks, while TinyStories and Wikitext were employed for language modeling experiments. QNLI dataset was used to evaluate downstream task perfomance. We implemented the distributed algorithm (DDP, Local SGD, DiLoCo and PALSGD) on three different distributed systems, referred to as Cluster A, Cluster B, and Cluster C which have varying GPU configurations and interconnects, as detailed below.
@@ B.1 Settings
  • VGG-16 on CIFAR-10: The CIFAR-10 dataset  [^7] https://www.cs.toronto.edu/~kriz/cifar.html is widely used in machine learning research, particularly for image recognition tasks. It contains 60,000 color images, each measuring 32×32 pixels, evenly distributed across ten distinct classes including airplanes, automobiles, birds, cats, deer, dogs, frogs, horses, ships, and trucks. The dataset consists of 50,000 training images and 10,000 test images, with each class represented by 6,000 images. We conducted experiments on CIFAR-10 using the VGG-16 architecture (Simonyan & Zisserman, 2014) to evaluate algorithm performance. Unlike prior experiments with smaller CNNs, this setup aims to leverage a deeper and more expressive model to assess the impact of algorithmic changes under realistic training conditions. The purpose of this experiment was not to measure training speed but to evaluate algorithm selection and perform ablation studies. The VGG-16 model we used follows the standard configuration defined in the PyTorch torchvision.models module, consisting of 13 convolutional layers followed by 3 fully connected layers. The convolutional layers use 3×3 filters with padding to preserve spatial resolution and are interleaved with ReLU activations and max-pooling layers for downsampling. After the convolutional blocks, the resulting feature maps are passed through an adaptive average pooling layer and flattened. The classifier consists of two fully connected layers with 4,096 units each (activated by ReLU and followed by dropout), and a final linear layer that maps to 10 output classes corresponding to the CIFAR-10 categories.

### v2

@@ 8 Discussion and Conclusion
  • We empirically validated PALSGD on image classification and language modeling tasks, demonstrating its effectiveness in reducing training time and improving model performance compared to baseline methods such as DDP, Local SGD and DiLoCo (Section 7, Figures 2, 4, and 3).
@@
Third, we do not include asynchronous baselines in this work. These methods have been studied extensively in Liu et al. (2024); Hadjis et al. (2016), and including them here would not only be redundant but could also obscure the focus of our study on semi-synchronous approaches. For this reason, we explicitly state their exclusion as a limitation here.
@@
Acknowledgement
Our deepest gratitude goes out to the anonymous reviewers and the Action Editor, whose invaluable insights substantially enhanced the quality of this manuscript. We also acknowledge that this research was enabled in part by computing resources, software, and technical assistance provided by Mila and the Digital Research Alliance of Canada. Man-Chung Yue is supported in part by Hong Kong Research Grants Council under the GRF project 15304422.
@@ A.1 Main Result
$\mathbb{E}[F(\hat{x}_{T})]-F(x^{*})\leq\tilde{O}\Bigg{(}\frac{LHR_{0}^{2}}{p}\cdot\exp(-\frac{pT}{\kappa H})+\frac{\sigma^{2}}{\mu KT}+\frac{\kappa H^{2}\sigma^{2}}{\mu T^{2}}\Bigg{)}.$
@@ A.2 Basic Definitions
For any $k\in[K],t=0,\cdots,T-1$, let $b_{k}^{(t)}$ denote a Bernoulli random variable with parameter $p$ (i.e., $b_{k}^{(t)}=1$ with probability $p$ and $b_{k}^{(t)}=0$ with probability 1p1-p). Let
@@ A.3 Main Technical Lemmas
$\leq$~ $12LH\cdot\sum_{s=t^{-}}^{t-1}\alpha_{s}^{2}\frac{1}{K}\sum_{k\in[K]}\mathbb{E}[F(x^{(s)})-F(x^{*})]+6H^{2}\alpha_{t^{-}}^{2}\sigma^{2}+\frac{p^{2}}{2H}\cdot\sum_{s=t^{-}}^{t-1}\frac{1}{K}\sum_{k\in[K]}\mathbb{E}[\|x_{k}^{(s)}-x^{(t^{-})}\|^{2}].$
@@ A.4 Proof of Theorem 2
$\leq$~ $(1-\mu\alpha_{t})\,\mathbb{E}\left[\|\bar{x}^{(t)}-x^{*}\|^{2}\right]+\left(\frac{24L\alpha_{t}^{2}}{K^{2}}-\frac{\alpha_{t}}{2}\right)\,\mathbb{E}\left[F(\bar{x}^{(t)})-F(x^{*})\right]$
@@
Next, by taking the weighted average from $t=0$ to T1T-1 with weight $w_{t}=(1-\mu\alpha)^{-(t+1)}$ and normalizing factor $W_{T}=\sum_{t=0}^{t-1}w_{t}$, we have
@@
$\leq$~ $\frac{1}{\alpha}\cdot\frac{1}{W_{T}}\sum_{t=0}^{T-1}\left(w_{t-1}d_{t}-w_{t}d_{t+1}\right)$
@@
$\leq$~ $\frac{1}{\alpha}\cdot\frac{1}{W_{T}}\sum_{t=0}^{T-1}\left(w_{t-1}d_{t}-w_{t}d_{t+1}\right)$
@@
$\leq$~ $\frac{1}{\alpha}\cdot\frac{1}{W_{T}}\left(d_{0}-w_{T}d_{t+1}\right)$
@@
$g(\alpha)=\frac{d_{0}}{\alpha}\exp(-\mu T\alpha)+81L\alpha^{2}H^{2}\sigma^{2}+\frac{12\alpha\sigma^{2}}{K}$
We wish to minimize $g(\alpha)$ with respect to $\alpha$. To do so, differentiate $g(\alpha)$ with respect to $\alpha$ we get
$g^{\prime}(\alpha)=-d_{0}\exp(-\mu T\alpha)\left(\frac{1}{\alpha^{2}}+\frac{\mu T}{\alpha}\right)+162LH^{2}\sigma^{2}\alpha+\frac{12\sigma^{2}}{K},$ (13)
and we want to find $\alpha^{*}$ such that $g^{\prime}(\alpha^{*})=0$. Since $g^{\prime\prime}(\alpha)>0$, such $\alpha^{*}$ is the minimizer of $g(\alpha)$.
However, it is complicated to get the accurate value of $\alpha^{*}$. But since we will ignore the poly-logarithmic terms in our final bound of $g(\alpha)$, it suffices to approximate $\alpha^{*}$ within poly-logarithmic factor. This would give a proper approximation of the the second and the third terms of $g(\alpha^{*})$. For the first term, we can use Equation equation 13 to derive the value of $\exp(-\mu T\alpha)$ and simplify the term to a polynomial of $\alpha$.
Suppose $\alpha^{*}=\frac{1}{\mu T}\ln A$. Our goal is to show that $A=\Theta(\operatorname{poly}(T,H,K,\sigma^{-1}))$. First, by plugging $\alpha^{*}=\frac{1}{\mu T}\ln A$ to equation 13 we get
$g^{\prime}(\alpha^{*})=-\frac{d_{0}}{A}\cdot\frac{(\ln A+1)\mu^{2}T^{2}}{\ln^{2}A}+\frac{162LH^{2}\sigma^{2}\ln A}{\mu T}+\frac{12\sigma^{2}}{K}=0.$ (14)
Simplifying both sides, we have
$\frac{d_{0}}{A}\cdot\frac{(\ln A+1)\mu^{2}T^{2}}{\ln^{2}A}=\frac{162LH^{2}\sigma^{2}\ln A}{\mu T}+\frac{12\sigma^{2}}{K},$
and
$\frac{\ln A+1}{A\ln^{2}A}=\frac{\sigma^{2}}{d_{0}\mu^{2}T^{2}}\cdot\left(\frac{162\kappa H^{2}\ln A}{T}+\frac{12}{K}\right).$
Assuming $T$ to be sufficiently large and $\frac{162\kappa H^{2}\ln A}{T}\ll\frac{12}{K}$, we can further simplify the above equality as
$\frac{\ln A+1}{A\ln^{2}A}=\frac{C\sigma^{2}}{d_{0}\mu^{2}T^{2}K}$ (15)
for some constant $C>12$. Notice that for any $A>4$, $\frac{\ln^{2}A}{\ln A+1}\leq A$ and $\frac{\ln^{2}A}{\ln A+1}\geq\frac{1}{\sqrt{A}}$, we have
$\frac{1}{\sqrt{A}}\leq\frac{\ln A+1}{A\ln^{2}A}\leq\frac{1}{A^{2}},$
for any $A\geq 4$. Plugging back to equation 15, we have that
$\frac{1}{\sqrt{A}}\leq\frac{C\sigma^{2}}{d_{0}\mu^{2}T^{2}K}\leq\frac{1}{A^{2}}.$
Therefore $A$ is polynomial in $T,H,K$ and $\sigma^{-1}$.
Now we upper-bound the minimal value of $g(\alpha)$. Since Theorem 2 also requires that $\alpha<\frac{p}{48LH}$, we need to discuss the following two cases:
If $\alpha^{*}<\frac{p}{48LH}$, then $g(\alpha)$ takes the minimum when $\alpha=\alpha^{*}$. Plugging in $\alpha^{*}=\frac{1}{\mu T}\ln A$ we have
$g(\alpha)$ $=g(\alpha^{*})$
$=\frac{d_{0}}{\alpha^{*}}\exp(-\mu T\alpha^{*})+81L(\alpha^{*})^{2}H^{2}\sigma^{2}+\frac{12\alpha^{*}\sigma^{2}}{K}$
$=\left(162LH^{2}\sigma^{2}\alpha^{*}+\frac{12\sigma^{2}}{K}\right)\cdot\frac{1}{\mu T}+81LH^{2}\sigma^{2}(\alpha^{*})^{2}+\frac{12\sigma^{2}\alpha^{*}}{K}$
$=\frac{162LH^{2}\sigma^{2}\ln A}{\mu^{2}T^{2}}+\frac{12\sigma^{2}}{\mu KT}+\frac{81LH^{2}\sigma^{2}\ln^{2}A}{\mu^{2}T^{2}}+\frac{12\sigma^{2}\ln A}{\mu KT}$
$=\tilde{O}\left(\frac{\kappa H^{2}\sigma^{2}}{\mu T^{2}}+\frac{\sigma^{2}}{\mu KT}\right).$
Here the step comes from equation 13, the third step come from $\alpha^{*}=\frac{1}{\mu T}\ln A$, and the last step comes from $A$ is polynomial in $T,H,K$ and $\sigma^{-1}$.
If $\alpha^{*}\geq\frac{p}{48LH}$, then since $g(\alpha)$ is convex in $\alpha$, it is monotonically non-increasing when $\alpha\leq\frac{p}{48LH}$. Thus $g(\alpha)$ takes the minimal value when $\alpha=\frac{p}{48LH}$. Therefore
$g(\alpha)=g(\frac{p}{48LH})\leq\tilde{O}(\frac{d_{0}LH}{p}\exp(-\frac{\mu pT}{48LH})+\frac{\kappa H^{2}\sigma^{2}}{\mu T^{2}}+\frac{\sigma^{2}}{\mu KT})$
Here we use $\frac{p}{48LH}\leq\alpha^{*}\leq\tilde{O}(\frac{1}{\mu T})$ on the second and the third terms in the second step.
Combining the two cases, we have
$\frac{1}{5}\cdot\frac{1}{W_{T}}\sum_{t=0}^{T-1}w_{t}e_{t}\leq{O}\left(\frac{d_{0}LH}{p}\exp(-\frac{\mu pT}{48LH})+\frac{\kappa H^{2}\sigma^{2}}{\mu T^{2}}+\frac{\sigma^{2}}{\mu KT}\right)$
@@ A.5.1 Proof of Lemma 2
where $b_{k}^{(t)}$ is a Bernoulli random variable that equals to $1$ with probability $p$ and 0 with probability 1p1-p. Also $\xi_{k}^{(t)}$ denotes the data sampled by the $k$-th server at step $t$. Let $\mathcal{F}_{t}=\{\xi_{k}^{(s)}\}_{s=0,\cdots,t-1,k\in[K]}\cup\{b_{k}^{(s)}\}_{s=0,\cdots,t-1,k\in[K]}$. Conditional on any fixed $\mathcal{F}_{t}$, we then rewrite each of the summands as
@@
$\leq$~ $\mathbb{E}_{\xi^{(t)}_{k}}[\|\nabla f(x_{k}^{(t)},\xi_{k}^{(t)})\|^{2}]$
$\leq$~ $\mathbb{E}_{\xi^{(t)}_{k}}[\|\nabla f(x_{k}^{(t)},\xi_{k}^{(t)})-\nabla f(x^{*},\xi_{k}^{(t)})+\nabla f(x^{*},\xi_{k}^{(t)})\|^{2}]$
$\leq$~ $\mathbb{E}_{\xi^{(t)}_{k}}[\|\nabla f(x_{k}^{(t)},\xi_{k}^{(t)})-\nabla f_{\xi^{(t)}_{k}}(x^{(t)})+\nabla f_{\xi^{(t)}_{k}}(x^{(t)})-\nabla f(x^{*},\xi_{k}^{(t)})+\nabla f(x^{*},\xi_{k}^{(t)})\|^{2}]$
$\leq$~ $3\cdot\mathbb{E}_{\xi^{(t)}_{k}}[\|\nabla f(x_{k}^{(t)},\xi_{k}^{(t)})-\nabla f_{\xi^{(t)}_{k}}(x^{(t)})\|^{2}+\|\nabla f_{\xi^{(t)}_{k}}(x^{(t)})-\nabla f(x^{*},\xi_{k}^{(t)})\|^{2}+\|\nabla f(x^{*},\xi_{k}^{(t)})\|^{2}]$
$\leq$~ $3L^{2}\mathbb{E}[\|x^{(t)}_{k}-x^{(t)}\|^{2}]+6L\mathbb{E}_{\xi^{(t)}_{k}}[f(x^{(t)},\xi^{(t)}_{k})-f(x^{*},\xi^{(t)}_{k})]+3\sigma^{2}$
$=$~ $3L^{2}\mathbb{E}[\|x^{(t)}_{k}-x^{(t)}\|^{2}]+6L\mathbb{E}[F(x^{(t)})-F(x^{*})]+3\sigma^{2}$ (17)
@@ A.5.2 Proof of Lemma 3
$\leq$~ $\frac{H}{K}\cdot\sum_{k\in[K]}\sum_{s=t^{-}}^{t-1}\mathbb{E}\left[\|x^{(t+1)}_{k}-x_{t}^{k}\|^{2}\right]$
$=$~ $\frac{H}{K}\cdot\sum_{k\in[K]}\sum_{s=t^{-}}^{t-1}\left(\frac{\alpha_{s}^{2}}{(1-p)}\cdot\mathbb{E}_{\xi_{k}^{(s)}}[\|\nabla f(x_{k}^{(s)},\xi_{k}^{(s)})\|^{2}]+\frac{\alpha_{s}^{2}\eta_{s}^{2}}{p}\cdot\mathbb{E}[\|x_{k}^{(s)}-x^{(t^{-})}\|^{2}]\right)$
$\leq$~ $\frac{H}{K}\cdot\sum_{k\in[K]}\sum_{s=t^{-}}^{t-1}\left(2\alpha_{s}^{2}\cdot\big{(}6L\mathbb{E}[F(x^{(s)})-F(x^{*})]+3\sigma^{2}\big{)}+\big{(}\frac{\alpha_{s}^{2}\eta_{s}^{2}}{p}+6\alpha_{s}^{2}L^{2}\big{)}\cdot\mathbb{E}[\|x_{k}^{(s)}-x^{(t^{-})}\|^{2}]\right)$
$\leq$~ $\frac{H}{K}\cdot\sum_{k\in[K]}\sum_{s=t^{-}}^{t-1}\left(6\alpha_{s}^{2}\cdot\big{(}2L\mathbb{E}[F(x^{(s)})-F(x^{*})]+\sigma^{2}\big{)}+\frac{p^{2}}{2H^{2}}\cdot\mathbb{E}[\|x_{k}^{(s)}-x^{(t^{-})}\|^{2}]\right)d$
$\leq$~ $12LH\cdot\sum_{s=t^{-}}^{t-1}\alpha_{s}^{2}\frac{1}{K}\sum_{k\in[K]}\mathbb{E}[F(x^{(s)})-F(x^{*})]+6H^{2}\alpha_{t^{-}}^{2}\sigma^{2}+\frac{p^{2}}{2H}\cdot\sum_{s=t^{-}}^{t-1}\frac{1}{K}\sum_{k\in[K]}\mathbb{E}[\|x_{k}^{(s)}-x^{(t^{-})}\|^{2}].$ (19)
@@ A.5.3 Proof of Lemma 4
$\leq$~ $\frac{p}{2H}\cdot\sum_{s=t^{-}}^{t-1}\left(\frac{p}{2H}\cdot\sum_{r=t^{-}}^{s-1}\Xi_{r}+6H\alpha^{2}\sum_{r=t^{-}}^{s-1}\sigma^{2}+12LH\cdot\sum_{r=t^{-}}^{s-1}\alpha^{2}e_{r}\right)$
@@
$\leq$~ $\frac{p^{2}}{4H}\cdot\sum_{s=t^{-}}^{t-2}\Xi_{s}+\sum_{s=t^{-}}^{t-2}(3Hp\alpha^{2}+6H\alpha^{2})\sigma^{2}+6H\alpha^{2}\sigma^{2}+\sum_{s=t^{-}}^{t-2}(6LHp\alpha^{2}+12LH\alpha^{2})e_{s}+12LH\alpha^{2}e_{t-1}.$
@@
$\leq$~ $9H^{2}\alpha^{2}\sum_{t=0}^{T}w_{t}\sigma^{2}+18LH\alpha^{2}\sum_{t=0}^{T}e_{t}\sum_{s=t^{-}}^{t-1}w_{s}$
$\leq$~ $9H^{2}\alpha^{2}\sum_{t=0}^{T}w_{t}\sigma^{2}+18LH\alpha^{2}\sum_{t=0}^{T}w_{t}e_{t}\sum_{s=t^{-}}^{t-1}(1+\frac{p}{H})^{s-t^{-}}$
$\leq$~ $9H^{2}\alpha^{2}\sum_{t=0}^{T}w_{t}\sigma^{2}+\frac{p}{16L}\sum_{t=0}^{T}w_{t}e_{t}$
@@ Appendix B Experimental Details
We conducted our experiments on three datasets: CIFAR-10, ImageNet-1K, and TinyStories. The CIFAR-10 and ImageNet-1K datasets were used for image classification tasks, while TinyStories was employed for language modeling experiments. We implemented the distributed algorithm (DDP, Local SGD, DiLoCo and PALSGD) on three different distributed systems, referred to as Cluster A, Cluster B, and Cluster C which have varying GPU configurations and interconnects, as detailed below.
@@ B.1 Settings
  • Small CNN on CIFAR-10 (Preliminary Simulation Experiments): The CIFAR-10 dataset  [^3] https://www.cs.toronto.edu/~kriz/cifar.html is widely used in machine learning research, particularly for image recognition tasks. It contains 60,000 color images, each measuring 32×32 pixels, evenly distributed across ten distinct classes including airplanes, automobiles, birds, cats, deer, dogs, frogs, horses, ships, and trucks. The dataset consists of 50,000 training images and 10,000 test images, with each class represented by 6,000 images. We employed a small CNN model in Table B.1 and the CIFAR10 dataset. We investigate accuracy degradation across varying numbers of workers, including extremely large worker configurations. Due to limited computing resources typical of academic institutions, experiments were performed on a simulated multi-worker environment using a single GPU and a relatively small-scale model. When controlling seeds and other conditions, such simulated experiments accurately represent actual distributed learning in terms of accuracy. However, execution time results are not comparable, as parallelization within a GPU differs fundamentally from distributed settings. Therefore, only accuracy results are reported.
  • VGG-16 on CIFAR-10: We conducted experiments on CIFAR-10 using the VGG-16 architecture (Simonyan & Zisserman, 2014) to evaluate algorithm performance. Unlike prior experiments with smaller CNNs, this setup aims to leverage a deeper and more expressive model to assess the impact of algorithmic changes under realistic training conditions. The VGG-16 model we used follows the standard configuration defined in the PyTorch torchvision.models module, consisting of 13 convolutional layers followed by 3 fully connected layers. The convolutional layers use 3×3 filters with padding to preserve spatial resolution and are interleaved with ReLU activations and max-pooling layers for downsampling. After the convolutional blocks, the resulting feature maps are passed through an adaptive average pooling layer and flattened. The classifier consists of two fully connected layers with 4,096 units each (activated by ReLU and followed by dropout), and a final linear layer that maps to 10 output classes corresponding to the CIFAR-10 categories.
@@
Table B.1: Architecture of the small CNN model used for simulation experiments
Layer Type Configuration
1 Convolution channels: 3 $\rightarrow$ 32, kernel size: 5, padding: 2
2 Activation ReLU
3 Max Pooling kernel size: 2, stride: 2
4 Convolution channels: 32 $\rightarrow$ 64, kernel size: 5, padding: 2
5 Activation ReLU
6 Max Pooling kernel size: 2, stride: 2
[…]
8 Activation ReLU
9 Fully Connected input: 1024, output: 10